In a reaction between Mg and hydrochloric acid:

Mg(s) + HCl(aq) —> MgCl2(aq) + H2(g)


Given the data below. Examine the data. What is the average rate of reaction?


Time


2

6

8


Mass of Magnesium

0.012

0.010

0.006

0.004


1
Expert's answer
2022-05-23T02:05:03-0400

The rate of a reaction ( V ) is the change in the concentration of a substance that reacts or is formed during this unit of time:

V = ∆C / ∆t

∆C = C2 - C1

∆t = t2 - t1

C2 — initial concentration

C1 — final concentration

t2 — initial time

t1 — final time

V = (0,012-0,004) / (8-0)= 0,008 / 8 = 0,001

ANSWER: 0,001